# Reducing GPU Memory Fragmentation via Spatio-Temporal Planning for Efficient Large-Scale Model Training

### v1

@@ 7. Implementation
STWeaver is implemented for PyTorch using about 5700 lines of Python and 3700 lines of C++. The plan synthesizer is implemented as a standalone tool, while the profiler and allocator are implemented as PyTorch’s PluggableAllocator (PyTorch, 2025), which can be loaded before training to take over the malloc and free API calls. This means that STWeaver is compatible with any PyTorch version and GPU platform that supports the PluggableAllocator interface. To capture temporal and spatial characteristics, STWeaver employs monkey patching for lightweight instrumentation, requiring no more than five lines of code in the original training framework.
@@
Runtime Allocator. At runtime, the allocator performs memory allocation according to the allocation plan. During training initialization, STWeaver uses native GPU memory allocation APIs to preallocate a contiguous memory block equal in size to the static memory pool and also initializes a caching allocator as a fallback. The runtime allocator then assigns address ranges within the static memory pool without issuing additional GPU memory API calls, thereby avoiding extra runtime overhead. To identify the current model layer during execution and route memory requests to the appropriate allocator, STWeaver leverages PyTorch’s hook APIs to track the execution of model modules. When a memory request arrives at runtime, the Request Matcher in the Runtime Allocator uses the current module information to determine whether the request should be handled by the static allocator according to the allocation plan, or by the dynamic allocator for online allocation.
8. Evaluation
To gain an in-depth understanding of STWeaver, we focus on the following aspects in the evaluation. (1) Performance. We show that STWeaver can reduce fragmentation memory by 79.2% on average (up to 100%), saving up to 56.3GB GPU memory across dense/sparse models trained with a variety of frameworks, configurations, and scales. (2) Overhead. We demonstrate that STWeaver’s impact on end-to-end training throughput is negligible in all cases, and our plan synthesizer can efficiently produce an allocation plan in minutes even under complex allocation requests. (3) Performance Breakdown. We study the individual performance of the static and dynamic allocators, and show their impacts on the final performance of STWeaver.
8.1. Experimental Setup
Testbed. STWeaver is evaluated on both NVIDIA and AMD GPU platforms. One configuration consists of 1 node equipped with an Intel Xeon Platinum 8358 128-Core CPU and 8 NVIDIA A800-80GB GPUs, which is used to evaluate various training optimization setups. The other has up to 16 nodes, each equipped with an Intel Xeon Platinum 8558 192-Core CPU and 8 NVIDIA H200-141GB GPUs, and is used for scalability evaluation. The AMD GPU platform has 8 nodes, each of which is equipped with AMD EPYC 7K62 48-Core Processor and 8 AMD MI210-64GB GPUs.
Models. We evaluate STWeaver on 7 representative large-scale dense and sparse Mixture-of-Expert (MoE) models. For dense models, we choose GPT-2 (Radford et al., 2019) and Llama2-7B (Touvron et al., 2023) for experiments on multiple training configurations. We use four models of varying sizes (including 7B, 14B, 32B, 72B) from the Qwen2.5 (Yang et al., 2024) series to demonstrate the scalability of our approach with respect to both model size and cluster size. For sparse models, we choose Qwen1.5-MoE-A2.7B (Team, 2024), a MoE model with 16 billion parameters to evaluate the efficiency of STWeaver on both multiple configurations as well as scalability and extendability on AMD platform.
Training Setup. We evaluate STWeaver with multiple training setups, in terms of training frameworks and training optimization techniques. For training frameworks, we choose the popular Megatron-LM (Shoeybi et al., 2019), Megatron-DeepSpeed (Smith et al., 2022), and Colossal-AI (Li et al., 2023). For training optimizations, we choose the pipeline parallelism schedule of Pipedream-1F1B (Narayanan et al., 2019), Virtual Pipeline (Narayanan et al., 2021b) as parallelism-based optimizations. For non-parallelism-based optimizations, we consider activation recomputation (Chen et al., 2016), offloading (Ren et al., 2021), and distributed optimizer (ZeRO (Rajbhandari et al., 2020)), which contains all kinds of memory optimizations (Duan et al., 2024).
Baselines. We compare STWeaver with state-of-the-art baselines, including:
  • PyTorch (Paszke et al., 2019). PyTorch employs a caching memory allocator for GPU memory management. It reduces the overhead of frequent native GPU API calls by reusing previously freed memory blocks, improving performance and memory efficiency.
  • PyTorch expandable_segments (PyTorch ES) (PyTorch, 2025). The expandable_segments allocator in PyTorch introduces support for virtual memory, allowing memory segments to grow dynamically as needed. This feature is only available in PyTorch versions 2.1 and above.
  • GMLake (Guo et al., 2024). GMLake leverages virtual memory stitching to unify non-contiguous memory blocks into a single virtual space for defragmenetation. We deployed it using the official Docker image provided in its repository (Group, 2023), whose PyTorch version is 2.0.
Metrics. We evaluate the performance of STWeaver using three key metrics. First, memory efficiency is the ratio of the max allocated memory to the max reserved memory as explained in §2.2. Building on this, the fragmentation ratio represents the proportion of reserved memory that is not actually utilized, which equals to (1 - memory efficiency). To measure the end-to-end throughput of training and evaluate the overhead of STWeaver, we choose FLOPS (floating point operations per second) as the throughput metric, which is calculated by training frameworks per training iteration.
8.2. Memory Efficiency and Defragmentation
@@
Models and Optimization Techniques. For the model and optimization combination test, the micro-batch sizes are set to the maximum feasible size that will not cause OOM following common practices (Narayanan et al., 2021b), i.e., 128, 4, and 8 for GPT-2, Llama2-7B, and Qwen1.5-MoE-A2.7B, respectively.
Our experiments used two training frameworks: Megatron-LM and Megatron-DeepSpeed. Megatron-LM is employed for configurations with pipeline parallelism ($PP$); for dense models, we set tensor parallelism ($TP$) to 4 and $PP$ to 2, while MoE models used $TP=4$, $PP=2$, and expert parallelism $EP=4$ to hold the model. In contrast, Megatron-DeepSpeed is used to evaluate techniques of ZeRO and ZeRO Offload. Due to architectural incompatibility of the Qwen1.5-MoE-A2.7B model with Megatron-DeepSpeed for ZeRO-like evaluations, we instead use Megatron-LM’s distributed optimizer and optimizer offload features as a substitute. Finally, since evaluating ZeRO and ZeRO Offload requires $DP>1$, the settings on GPT-2 and Qwen1.5-MoE-A2.7B model exceed the eight A800 GPU memory capacity, we tested ZeRO with recomputation as an alternative for these cases.
@@
Figure 7 shows the memory efficiency comparison. We can observe that for dense models that do not have dynamic layers, STWeaver can achieve $>$95% (up to 100%) memory efficiency (i.e., fragmentation ratio ¡ 5%) in all cases, demonstrating the effectiveness of our spatio-temporal planning mechanism. In comparison, PyTorch 2.3 produces 57.2% to 93.8% memory efficiency, GMLake produces 53.6% to 96.5% memory efficiency, and PyTorch ES yields 62.4% to 94.9% memory efficiency. Compared to the baselines, STWeaver reduces fragmentation memory by 88.1%, 77.1%, and 76.0% on average, up to 100%, reducing reserved memory up to 11GB (i.e., 13.8% of GPU memory). The most significant fragmentation reduction appears in the case of GPT-2 with ZeRO-3 and recomputation, which is because the weight size of GPT-2 is relatively small compared to the other 2 models, and thus the proportion of activation tensors (whose lifecycle is affected by recomputation) among all the tensors is considerably larger than that of Llama2-7B.
@@
For the MoE model with dynamic layers, STWeaver still shows 93.8% to 97.8% memory efficiency in the evaluated cases, reducing the fragmentation ratio to 4.3% on average. Compared to PyTorch, GMLake, and PyTorch ES, whose fragmentation ratios are 17.7%, 20.3%, and 6.9%, respectively. STWeaver occurs less fragmentation memory of 77.1%, 78.3%, and 39.4%, respectively. In the MoE test, tuning the default GMLake defragmentation threshold (fragLimit) from 512 MB to 64 MB increased memory efficiency to 97.73% but reduced training performance by 56.4% over 50 iterations. The 64 MB threshold caused unstable virtual memory pools under MoE’s dynamic allocations, leading to frequent virtual memory operations (up to 1500 times per iteration, each taking around 30ms). A 512 MB threshold optimally balances memory efficiency and training performance.
@@
Training Scales. We demonstrate the scalability of STWeaver on the two different GPU platforms. On the AMD platform, we train the Llama2-7B and Qwen1.5-MoE-A2.7B models on 4 nodes (32 GPUs) and 8 nodes (64 GPUs), respectively. We excluded GMLake and PyTorch ES from this study, as GMLake does not support AMD GPUs, and the features of PyTorch ES are unavailable in our platform’s PyTorch version (2.0). All the training experiments are conducted with recomputation. As shown in Figure 8(a), STWeaver scales well for both the dense and MoE models. The memory efficiency on both models achieves over 90%, and up to 99.7%. In contrast, the PyTorch caching allocator exhibits memory efficiency below 60% across all scales of the Llama2-7B model. This result shows that STWeaver can reduce fragmentation memory of 22.8 GB, which is 35.6% of GPU memory. Moreover, for MoE models, when the cluster size increases from 32 to 64 GPUs, the memory efficiency drops below 80%.
@@
To further investigate the scalability of STWeaver’s memory efficiency as model and cluster sizes are concurrently augmented, we use four models of varying sizes from the Qwen2.5 series, including 7B, 14B, 32B, and 72B, on 8 to 128 NVIDIA H200 GPUs. The training configurations are either recomputation, as a memory optimization technique, or virtual pipeline, as a parallelism optimization strategy, thereby demonstrating STWeaver’s scalability across diverse scenarios. GMLake is not included since it does not support PyTorch 2.6 on the current platform.
@@
Under recomputation settings (Figure 8(b)), STWeaver achieves 99.1% memory efficiency, reducing fragmentation by over 98.5% and 98.4% compared to PyTorch 2.6 and PyTorch ES, respectively, saving 37.9 GB GPU memory on average, up to 56.3 GB. PyTorch ES showed throughput degradation: 15.0% lower than PyTorch for the 32B model on 32 GPUs, while STWeaver’s throughput matched PyTorch within 0.02%. For the 72B model on 64 GPUs, PyTorch faced OOM errors due to fragmentation, and PyTorch ES was 20.1% slower than STWeaver. PyTorch ES’s overhead stems from frequent virtual memory API calls, whereas STWeaver maintains high efficiency with minimal runtime penalties.
@@
Under virtual pipeline settings as shown in Figure 8(c), STWeaver achieves memory efficiency over 99% in all cases; reduce fragmentation memory over 97.6% and 97.4% compared to PyTorch 2.6 and PyTorch ES, respectively, saving GPU memory of 15.7 GB on average. We find that with the scaling of model and cluster sizes, the memory efficiency of PyTorch and PyTorch ES declined by 10.9% and 15.0%, respectively, while STWeaver differs within 0.7%.
@@
When training the 14B model on 16 GPUs, only STWeaver successfully completes the training without out-of-memory (OOM) error by reducing fragmentation. To avoid OOM, PyTorch and PyTorch ES require disabling virtual pipeline, increasing the tensor parallelism degree, or introducing recomputation. The original training configuration outperforms these adjustments in training throughput by 5.4% to 32.5%, as shown in Table 1. This indicates that by reducing fragmentation, STWeaver enables more efficient training configurations and yields performance improvements.
@@
Config PyTorch PyTorch ES STWeaver Throughput (TFLOPS)
@@
$TP=4$checkmark $\checkmark$ $\checkmark$ 431.5
Micro-Batch Sizes. Given that activation memory usage during training is directly proportional to microbatch size, and that larger microbatch sizes typically enhance operator computational efficiency (Narayanan et al., 2021b), we conducted further experiments across a range of microbatch sizes. We conduct the experiments for micro-batch sizes 1, 2, 4, 8, 16, 32, and 64, training Llama2-7B with recomputation on Megatron-LM. As shown in 9, STWeaver yields the best and similar (around 99%) memory efficiency regardless of the micro-batch size, while the other allocators generally performs worse as the micro-batch size increases, mostly because the increasing size of the activation tensors affected by recomputation. This proves STWeaver’s robustness against memory-related training configurations in practice.
Training Frameworks. To evaluate STWeaver’s generalizability across high-level training frameworks, we also apply STWeaver to Colossal-AI (Li et al., 2023), another representative training framework shipped with a variety of memory optimizations. We train GPT-2 on Colossal-AI with tensor offload and ZeRO-3 (Rajbhandari et al., 2020) with two different batch sizes. As depicted in Figure 12, STWeaver still performs better than the other allocators, demonstrating STWeaver’s general applicability across training frameworks.
@@ 8.3. Overhead Analysis
We next evaluate STWeaver’s potential impact on the end-to-end training throughput, as well as the efficiency of the allocation profiler and plan synthesizer facing different numbers of allocation requests.
Overhead of Allocators in Training Throughput. Figure 12 shows the normalized end-to-end training throughput when training the 3 test models on Megatron-LM using different allocators. Specifically, GMLake is normalized against PyTorch 2.0, while PyTorch ES and STWeaver are normalized against PyTorch 2.3 for fairness. All the experiment settings adopt recomputation. We can see that none of the allocators incur noticeable throughput degradation. In particular, STWeaver’s throughput difference with the vanilla PyTorch 2.3 is $<$0.05% in all cases, which are most likely due to hardware performance fluctuation. It is worth noting that virtual memory–based GPU memory allocation methods have shown significant drops in training throughput under specific scenarios as discussed in §8.2. GMLake exhibits such behavior in MoE models, and PyTorch ES demonstrates it in recomputation-heavy settings. While these approaches help reduce memory fragmentation, the runtime overhead introduced by virtual memory operations can become non-negligible, ultimately impacting training performance.
The above throughput comparison uses identical training configurations. Thanks to STWeaver’s ability to reduce GPU memory usage without incurring extra runtime overhead, it enables the use of more memory-intensive configurations without triggering out-of-memory (OOM) errors. As a result, STWeaver can achieve higher training throughput.
@@
Table 2. Profile and plan synthesis time in different training configuration. $Num$ is the number of requests within one iteration. -N and -R represent the configuration without/with recomputation, respectively.
Config $Num$ $T_{profile}(s)$ $T_{plan}(s)$

### v2

@@ 8. Implementation
STAlloc is implemented for PyTorch using about 3100 lines of Python and 2300 lines of C/C++. The plan synthesizer is implemented as a standalone tool, while the profiler and allocator are implemented as PyTorch’s PluggableAllocator (pytorch-doc), which can be loaded before training to take over the malloc and free API calls. This means that STAlloc is compatible with any PyTorch version and GPU platform that supports the PluggableAllocator interface. To capture temporal and spatial characteristics, STAlloc employs monkey patching for lightweight instrumentation, requiring no more than five lines of code in the original training framework.
@@
Runtime Allocator. At runtime, the allocator performs memory allocation according to the allocation plan. During training initialization, STAlloc uses native GPU memory allocation APIs to preallocate a contiguous memory block equal in size to the static memory pool and also initializes a caching allocator as a fallback. The runtime allocator then assigns address ranges within the static memory pool without issuing additional GPU memory API calls, thereby avoiding extra runtime overhead. To identify the current model layer during execution and route memory requests to the appropriate allocator, STAlloc leverages PyTorch’s hook APIs to track the execution of model modules. When a memory request arrives at runtime, the Request Matcher in the Runtime Allocator uses the current module information to determine whether the request should be handled by the static allocator according to the allocation plan, or by the dynamic allocator for online allocation.
@@
9. Evaluation
To gain an in-depth understanding of STAlloc, we focus on the following aspects in the evaluation. (1) Performance. We show that STAlloc can reduce fragmentation memory by 85.1% on average (up to 100%), saving up to 56.3GB GPU memory across dense/sparse models trained with a variety of frameworks, configurations, and scales. (2) Overhead. We demonstrate that STAlloc’s impact on end-to-end training throughput is negligible in all cases, and our plan synthesizer can efficiently produce an allocation plan in minutes even under complex allocation requests. (3) Performance Breakdown. We study the individual performance of the static and dynamic allocators, and show their impacts on the final performance of STAlloc.
9.1. Experimental Setup
Testbed. STAlloc is evaluated on both NVIDIA and AMD GPU platforms. One configuration consists of 1 node equipped with an Intel Xeon Platinum 8358 128-Core CPU and 8 NVIDIA A800-80GB GPUs, which is used to evaluate various training optimization setups. The other has up to 16 nodes, each equipped with an Intel Xeon Platinum 8558 192-Core CPU and 8 NVIDIA H200-141GB GPUs, and is used for scalability evaluation. The AMD GPU platform has 8 nodes, each of which is equipped with AMD EPYC 7K62 48-Core Processor and 8 AMD MI210-64GB GPUs.
Models. We evaluate STAlloc on 7 representative large-scale dense and sparse Mixture-of-Expert (MoE) models. For dense models, we choose GPT-2 (openai-gpt2) and Llama2-7B (llama) for experiments on multiple training configurations. We use four models of varying sizes (including 7B, 14B, 32B, 72B) from the Qwen2.5 (qwen2.5) series to demonstrate the scalability of our approach with respect to both model size and cluster size. For sparse models, we choose Qwen1.5-MoE-A2.7B (qwen1.5-moe), a MoE model with 16 billion parameters to evaluate the efficiency of STAlloc on both multiple configurations as well as scalability and extendability on AMD platform.
Training Setup. We evaluate STAlloc with multiple training setups, in terms of training frameworks and training optimization techniques. For training frameworks, we choose the popular Megatron-LM (megatron), and Colossal-AI (colossal-ai). For training optimizations, we choose the pipeline parallelism schedule of Pipedream-1F1B (pipedream-1f1b), Virtual Pipeline (megatron2) as parallelism-based optimizations. For non-parallelism-based optimizations, we consider activation recomputation (recompute1), offloading (zero-offload), and distributed optimizer (ZeRO (zero)), which contains all kinds of memory optimizations (training-survey).
Baselines. We compare STAlloc with state-of-the-art baselines, including:
  • PyTorch (pytorch). PyTorch employs a caching memory allocator for GPU memory management. It reduces the overhead of frequent native GPU API calls by reusing previously freed memory blocks, improving performance and memory efficiency.
  • PyTorch expandable_segments (PyTorch ES) (pytorch-doc). The expandable_segments allocator in PyTorch introduces support for virtual memory, allowing memory segments to grow dynamically as needed. This feature is only available in PyTorch versions 2.1 and above.
  • GMLake (gmlake). GMLake leverages virtual memory stitching to unify non-contiguous memory blocks into a single virtual space for defragmenetation. We deployed it using the official Docker image provided in its repository (glake_repo), whose PyTorch version is 2.0.
Metrics. We evaluate the performance of STAlloc using three key metrics. First, memory efficiency is the ratio of the max allocated memory to the max reserved memory as explained in §2.2. Building on this, the fragmentation ratio represents the proportion of reserved memory that is not actually utilized, which equals to (1 - memory efficiency). To measure the end-to-end throughput of training and evaluate the overhead of STAlloc, we choose FLOPS (floating point operations per second) as the throughput metric, which is calculated by training frameworks per training iteration.
9.2. Memory Efficiency and Defragmentation
@@
Models and Optimization Techniques. For the model and optimization combination test, the micro-batch sizes are set to the maximum feasible size that will not cause OOM following common practices (megatron2), i.e., 128, 4, and 8 for GPT-2, Llama2-7B, and Qwen1.5-MoE-A2.7B, respectively.
@@
Our experiments used Megatron-LM as the training framework. Figure 8 shows the comparison of memory efficiency. We can observe that for dense models that do not have dynamic layers, STAlloc can achieve >>95% (up to 100%) memory efficiency (i.e., fragmentation ratio ¡ 5%) in all cases, demonstrating the effectiveness of our spatio-temporal planning mechanism. In comparison, PyTorch 2.3 produces 57.1% to 90.6% memory efficiency, GMLake produces 45.1% to 88.1% memory efficiency, and PyTorch ES yields 62.4% to 93.2% memory efficiency. Compared to the baselines, STAlloc reduces fragmentation memory by 90.3%, 93.4%, and 87.8% on average, up to 100%, reducing reserved memory up to 14.4GB (i.e., 18% of GPU memory). The most significant fragmentation reduction appears in the case of GPT-2 with ZeRO and recomputation, which is because the weight size of GPT-2 is relatively small compared to the other 2 models, and thus the proportion of activation tensors (whose lifecycle is affected by recomputation) among all the tensors is considerably larger than that of Llama2-7B.
@@
For the MoE model with dynamic layers, STAlloc still shows 93.7% to 97.8% memory efficiency in the evaluated cases, reducing the fragmentation ratio to 4.3% on average. Compared to PyTorch, GMLake, and PyTorch ES, whose fragmentation ratios are 17.7%, 20.3%, and 6.9%, respectively. STAlloc occurs less fragmentation memory of 74.9%, 77.2%, and 34.0%, respectively. In the MoE test, tuning the default GMLake defragmentation threshold (fragLimit) from 512 MB to 64 MB increased memory efficiency to 97.73% but reduced training performance by 56.4% over 50 iterations. The 64 MB threshold caused unstable virtual memory pools under MoE’s dynamic allocations, leading to frequent virtual memory operations (up to 1500 times per iteration, each taking around 30ms). A 512 MB threshold optimally balances memory efficiency and training performance.
@@
Training Scales. We demonstrate the scalability of STAlloc on the two different GPU platforms. On the AMD platform, we train the Llama2-7B and Qwen1.5-MoE-A2.7B models on 4 nodes (32 GPUs) and 8 nodes (64 GPUs), respectively. We excluded GMLake and PyTorch ES from this study, as GMLake does not support AMD GPUs, and the features of PyTorch ES are unavailable in our platform’s PyTorch version (2.0). All the training experiments are conducted with recomputation. As shown in Figure 9(a), STAlloc scales well for both the dense and MoE models. The memory efficiency on both models achieves over 90%, and up to 99.7%. In contrast, the PyTorch caching allocator exhibits memory efficiency below 60% across all scales of the Llama2-7B model. This result shows that STAlloc can reduce fragmentation memory of 22.8 GB, which is 35.6% of GPU memory. Moreover, for MoE models, when the cluster size increases from 32 to 64 GPUs, the memory efficiency drops below 80%.
@@
To further investigate the scalability of STAlloc’s memory efficiency as model and cluster sizes are concurrently augmented, we use four models of varying sizes from the Qwen2.5 series, including 7B, 14B, 32B, and 72B, on 8 to 128 NVIDIA H200 GPUs. The training configurations are either recomputation, as a memory optimization technique, or virtual pipeline, as a parallelism optimization strategy, thereby demonstrating STAlloc’s scalability across diverse scenarios. GMLake is not included since it does not support PyTorch 2.6 on the current platform.
@@
Under recomputation settings (Figure 9(b)), STAlloc achieves 99.1% memory efficiency, reducing fragmentation by over 98.5% and 98.4% compared to PyTorch 2.6 and PyTorch ES, respectively, saving 37.9 GB GPU memory on average, up to 56.3 GB. PyTorch ES showed throughput degradation: 15.0% lower than PyTorch for the 32B model on 32 GPUs, while STAlloc’s throughput matched PyTorch within 0.02%. For the 72B model on 64 GPUs, PyTorch faced OOM errors due to fragmentation, and PyTorch ES was 20.1% slower than STAlloc. PyTorch ES’s overhead stems from frequent virtual memory API calls, whereas STAlloc maintains high efficiency with minimal runtime penalties.
@@
Under virtual pipeline settings as shown in Figure 9(c), STAlloc achieves memory efficiency over 99% in all cases; reduce fragmentation memory over 97.6% and 97.4% compared to PyTorch 2.6 and PyTorch ES, respectively, saving GPU memory of 15.7 GB on average. We find that with the scaling of model and cluster sizes, the memory efficiency of PyTorch and PyTorch ES declined by 10.9% and 15.0%, respectively, while STAlloc differs within 0.7%.
@@
When training the 14B model on 16 GPUs, only STAlloc successfully completes the training without out-of-memory (OOM) error by reducing fragmentation. To avoid OOM, PyTorch and PyTorch ES require disabling virtual pipeline, increasing the tensor parallelism degree, or introducing recomputation. The original training configuration outperforms these adjustments in training throughput by 5.4% to 32.5%, as shown in Table 1. This indicates that by reducing fragmentation, STAlloc enables more efficient training configurations and yields performance improvements.
@@
Config PyTorch PyTorch ES STAlloc Throughput (TFLOPS)
@@
$TP=4$ $\checkmark$ $\checkmark$ $\checkmark$ 431.5
Micro-Batch Sizes. Given that activation memory usage during training is directly proportional to microbatch size, and that larger microbatch sizes typically enhance operator computational efficiency (megatron2), we conducted further experiments across a range of microbatch sizes. We conduct the experiments for micro-batch sizes 1, 2, 4, 8, 16, 32, and 64, training Llama2-7B with recomputation on Megatron-LM. As shown in Figure 10, STAlloc yields the best and similar (around 99%) memory efficiency regardless of the micro-batch size, while the other allocators generally performs worse as the micro-batch size increases, mostly because the increasing size of the activation tensors affected by recomputation. This proves STAlloc’s robustness against memory-related training configurations in practice.
Training Frameworks. To evaluate STAlloc’s generalizability across high-level training frameworks, we also apply STAlloc to Colossal-AI (colossal-ai), another representative training framework shipped with a variety of memory optimizations. We train GPT-2 on Colossal-AI with tensor offload and ZeRO-3 (zero) with two different batch sizes. As depicted in Figure 13, STAlloc still performs better than the other allocators, demonstrating STAlloc’s general applicability across training frameworks.
@@ 9.3. Overhead Analysis
We next evaluate STAlloc’s potential impact on the end-to-end training throughput, as well as the efficiency of the allocation profiler and plan synthesizer facing different numbers of allocation requests.
Overhead of Allocators in Training Throughput. Figure 13 shows the normalized end-to-end training throughput when training the 3 test models on Megatron-LM using different allocators. Specifically, GMLake is normalized against PyTorch 2.0, while PyTorch ES and STAlloc are normalized against PyTorch 2.3 for fairness. All the experiment settings adopt recomputation. We can see that none of the allocators incur noticeable throughput degradation. In particular, STAlloc’s throughput difference with the vanilla PyTorch 2.3 is <<0.05% in all cases, which are most likely due to hardware performance fluctuation. It is worth noting that virtual memory–based GPU memory allocation methods have shown significant drops in training throughput under specific scenarios as discussed in §9.2. GMLake exhibits such behavior in MoE models, and PyTorch ES demonstrates it in recomputation-heavy settings. While these approaches help reduce memory fragmentation, the runtime overhead introduced by virtual memory operations can become non-negligible, ultimately impacting training performance.
The above throughput comparison uses identical training configurations. Thanks to STAlloc’s ability to reduce GPU memory usage without incurring extra runtime overhead, it enables the use of more memory-intensive configurations without triggering out-of-memory (OOM) errors. As a result, STAlloc can achieve higher training throughput.
@@
Table 2. Profile and plan synthesis time in different training configuration. NumNum is the number of requests within one iteration. -N and -R represent the configuration without/with recomputation, respectively.
Config NumNum $T_{profile}(s)$ $T_{plan}(s)$